# LLM Knows Geometry Better than Algebra: Numerical Understanding of LLM-Based Agents in A Trading Arena

### v1

@@ 9 Ivy 3 100,000 Marketing Manager
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
Table 1: Agent Details. “Year” is the investment duration, and “Cash” is the initial capital.
ID Name DPS Historical Close Prices Quantity
1 A 22 454.17, $\dots$, 445.60 1,200
2 B 23 354.17, $\dots$, 465.80 1,000
3 C 25 500.47, $\dots$, 440.60 1,600
$\dots$ $\dots$ $\dots$ $\dots$ $\dots$
Table 2: Stock Details. “DPS” is the Dividend Per Share, and “Quantity” is the initial share count.
4 Experimental Results
4.1 Experimental Setup
@@
To evaluate the ability of LLMs to analyze and process data, we developed the Agent Trading Arena, a controlled environment that isolates external factors. The system’s workflow is illustrated in Figure 2. Within this environment, agents discuss stock market trends, analyze stock data, and engage in trading activities. Each agent can execute multiple trades per day and reflect on all trades at the end of each trading session.The Agent Trading Arena allows for adjustable numbers of agents and stocks; in our experiment, we deployed at least nine agents and three stocks. All agents were provided with the same initial capital to ensure identical starting conditions. Detailed information about each agent is presented in Table 1, and stock-specific data is available in Table 2. To further validate our findings, we selected a subset of NASDAQ Stock dataset for portfolio investment. For more details, refer to Appendix B.
Evaluation Metrics.
Each agent was assigned varying capital based on their roles, and performance was evaluated using the following metrics:
1) Total Return (TR): Measures the overall performance of the strategy, calculated as: $\mathrm{TR}={(C_{1}-C_{0})}/{C_{0}}$, where $C_{0}$ is the initial asset value and $C_{1}$ is the final asset value.
2) Win Rate (WR): Represents the proportion of winning trades, calculated as: $\mathrm{WR}={N_{w}}/{N_{t}}$, where $N_{w}$ is the number of winning trades and $N_{t}$ is the total number of trades.
@@
3) Sharpe Ratio (SR): Evaluates investment returns relative to risk, calculated as: $\mathrm{SR}={(R_{p}-R_{f})}/{\sigma_{p}}$, where $R_{p}$ is the mean daily return, $\sigma_{p}$ is the standard deviation of daily returns, and $R_{f}$ is the risk-free return, set to 0 as defined in SocioDojo (Cheng and Chin, 2024).
4) Mean Daily Return (Mean): Indicates the average return per day during the trading period.
@@
5) Standard Deviation of Daily Returns (Std): Reflects the volatility of daily returns, indicating the risk associated with the strategy.
4.2 Comparative Experiments
We conducted experiments to assess the real-time data analysis and reasoning capabilities of LLM-based agents, focusing on how textual and visual representations influence decision-making. First, we explored the impact of textual and visual representations in dynamic environments. Next, we incorporated the reflection module to enhance agents’ reasoning and data interpretation, examining how reflective reasoning influences decision-making. Finally, we validated the model’s effectiveness through stock investment simulations on NASDAQ Stock dataset, assessing agents’ adaptability and decision-making in real-world scenarios.
LLMs Textual Visual TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
w/o Reflection LLaMa-3 (Dubey et al., 2024) 4.0934 0.6857 2.2700 62.5001 0.1963
DeepSeek (Liu et al., 2024) 27.3078 2.2478 2.7553 92.2078 1.5303
Qwen-2.5 (Yang et al., 2024b) 30.3740 2.6795 1.4588 93.7500 1.7025
Gemini-1.5 (Reid et al., 2024) 14.3193 1.2210 2.7776 83.8384 1.7293
19.0389 1.6038 1.1654 90.9091 1.3761
23.9649 1.9809 1.3342 100.0000 1.4847
GPT-4o (Hurst et al., 2024) 13.0369 1.1386 1.9285 54.5455 0.5904
17.0661 1.4539 1.5040 72.7272 0.9668
26.1806 2.1574 2.0578 90.9091 1.0484
w/ Reflection LLaMa-3 (Dubey et al., 2024) 10.2458 1.6669 2.4034 66.6667 0.6936
DeepSeek (Liu et al., 2024) 30.6238 2.4696 1.5121 100.0000 1.6332
Qwen-2.5 (Yang et al., 2024b) 38.9113 3.3244 1.0726 100.0000 3.0994
Gemini-1.5 (Reid et al., 2024) 29.4511 2.5185 3.0116 93.5064 2.2100
37.0054 2.9111 1.2349 100.0000 2.3574
41.3264 3.1946 0.1588 100.0000 20.1128
GPT-4o (Hurst et al., 2024) 33.6508 2.6713 2.1711 98.7013 2.1417
35.7622 2.8206 0.6782 100.0000 4.1590
47.6851 3.6095 0.5327 100.0000 6.7765
Table 3: Performance Comparison Using Textual and Visual Data w/o and w/ Reflection. Evaluation of agent portfolios show that visual approaches outperformed textual ones, with the best results achieved by combining both. LLMs demonstrated a significant advantage in TR and SR, with visual inputs yielding greater improvements than textual inputs. The best and second-best results are highlighted in bold and underlined.
Trials with Textual or Visual Input.
To enhance LLMs’ understanding of complex data in the Agent Trading Arena, we transitioned from textual numerical inputs to visualized formats, including scatter plots, line charts, and bar charts. Three types of visualizations were used: daily K-line charts, transaction histories, and quantities traded by each agent. The experimental setup for the visualization group is shown in Figure 3 and detailed in Appendix C. In the Arena, LLMs without image input capabilities received only text input and did not perform reflection. For image-enabled LLMs, the first agent received only visual input, the second received both textual and visual input, and the others received only textual input. None of the agents had reflection capabilities. For details on the selected LLMs, please refer to Appendix E.
We conducted experiments across different LLMs and the results for Gemini-1.5 (Reid et al., 2024) and GPT-4o (Hurst et al., 2024) in Table 3 show that agents with visual numerical input outperformed those with textual input alone. Agents receiving both textual and visual input achieved the best performance. This combination enables agents to focus on local details while also understanding overall trends, leading to optimal performance. This suggests that LLMs exhibit better reasoning skills with visual geometric data rather than textual numerical data, highlighting a clear distinction in their strengths in geometry and algebra.
Strategy TR $\uparrow$ SR $\uparrow$
MACD (Chong and Ng, 2008) 7.18 0.173
StockFormer (Gao et al., 2023) 9.05 0.073
TimesNet (Wu et al., 2023) 11.74 0.203
GPT-4o + Textual 8.69 0.167
GPT-4o + Visual 9.91 0.195
GPT-4o + Textual + Visual (Ours) 12.23 0.291
Table 4: Reproduced Backtesting Performance on NASDAQ Stock Dataset. The agent with textual and visual input outperformed the NASDAQ-100 by 53.97% in SR during the same period.
LLMs T V TR $\uparrow$ SR $\uparrow$
LLaMa-3 100.00 100.00
DeepSeek + 35.03 + 45.63
Gemini-1.5 + 10.99 + 14.93
+ 11.49 + 29.47
GPT-4o + 10.39 + 28.29
+ 17.18 + 40.77
LLMs TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
LLaMa-3 (Dubey et al., 2024) 100.0000 100.0000 100.0000 100.0000 100.0000
Gemini-1.5 (Reid et al., 2024) + 3.8631 + 5.9249 + 12.0426 - 4.0010 + 2.2920
DeepSeek (Liu et al., 2024) + 3.1317 + 3.1309 + 1.7169 + 9.8039 + 1.7945
Qwen-2.5 (Yang et al., 2024b) + 7.4578 + 7.3907 - 2.4801 + 50.0000 + 8.0524
GPT-4o (Hurst et al., 2024) + 17.7636 + 16.9041 + 2.4136 + 37.9310 + 14.2775
Table 6: Performance Comparison in Trading Decisions w/ Reflection, Competing Pairwise with LLaMa-3 Using Textual Data. The results show that in competition with reflection using textual numerical data, GPT-4o and Qwen-2.5 outperformed other models.
Trials with Reflection and Textual or Visual Input.
Building upon the previous experiments, we incorporated the reflection module to investigate its impact on the experimental outcomes. By introducing this module, we aimed to determine whether reflective processing could improve strategies and adaptability in complex financial scenarios.
Our experiments with diverse LLMs demonstrate that reflective agents consistently outperformed non-reflective agents in stock trading. As shown in Table 3, text-only LLMs with reflection modules achieved enhanced returns, with GPT-4o (Hurst et al., 2024) reaching 33.65% total return using textual input alone, which increased to 47.70% with visual data. Combining both input types resulted in a 41.74% higher return compared to using textual input alone. Compared to agents without the reflection module, including the reflection module significantly enhanced each agent’s stock investment performance. LLMs with the reflection module showed the greatest improvement in agents with visual inputs, compared to those with textual inputs, in both the Total Return and Sharpe Ratio. This indicates that incorporating reflection further widens the gap between LLMs’ geometric reasoning with visual data and algebraic reasoning with textual numerical data, highlighting their strengths in both areas.
Simulation with NASDAQ Stock dataset.
To further validate our findings, we conducted a two-month investment simulation on NASDAQ Stock dataset, starting with an initial capital of 100,000 units for portfolio investment. NASDAQ Stock dataset from Yahoo Finance spans July 3, 2023, to October 29, 2024, excluding weekends and holidays. Detailed information about the models used is provided in Appendix B.
The results in Table 4 show that, despite StockFormer (Gao et al., 2023) and TimesNet (Wu et al., 2023) benefiting from longer training periods and larger datasets, they underperformed compared to our model. Notably, our model achieved superior results without additional training, relying solely on historical stock data for decision-making. The Sharpe ratios for Apple, NASDAQ-100, and S&P 500 during the same period were 0.097, 0.189, and 0.205, respectively. The agent with textual and visual input outperformed the NASDAQ-100 and S&P 500 in Sharpe Ratio by 53.97% and 41.95%, respectively. Moreover, only visual input outperformed textual input, further highlighting LLMs’ stronger geometric reasoning abilities.
4.3 Ablation Experiments
Impact of Modality on LLM Competitiveness.
We employed a relative evaluation method for this experiment. The first and second agents used various LLMs in textual and visual settings, respectively, while the remaining agents were based on LLaMa-3 (Dubey et al., 2024) as the baseline. This setup aimed to explore the impact of different agents and modalities on LLM performance. The results are shown in Table 5. The findings indicate that DeepSeek (Liu et al., 2024) exhibited stronger competitive performance across different LLM environments, suggesting unique strengths or optimizations that enable it to adapt more effectively to the task’s complexities.
Competition Among Different LLMs with Reflection.
To investigate the role of reflection in competition dynamics using textual numerical data between LLMs, we conducted an ablation study in the Agent Trading Arena. Using a relative evaluation approach, the first agent employed various LLMs, while eight LLaMa-3-based agents served as the baseline for a fair comparison. This setup effectively isolated the influence of each LLM on performance. As shown in Table 6, in competition with reflection, GPT-4o (Hurst et al., 2024) and Qwen-2.5 (Yang et al., 2024b) outperformed other models, consistent with the findings in Table 3.
@@ 5 Conclusion
In this paper, we introduced the Agent Trading Arena, a zero-sum game designed to simulate complex economic systems and evaluate LLMs on numerical reasoning tasks. Our results show that while LLMs struggle with plain-text numerical data (algebraic reasoning), their performance significantly improves when presented with visual data (geometric reasoning). This highlights the advantage of visual representations in supporting numerical reasoning, particularly in complex scenarios. The integration of a reflection module further enhances model performance, allowing LLMs to analyze better and interpret data. We validated these findings on NASDAQ Stock dataset, demonstrating that LLMs excel in visual geometric reasoning tasks, suggesting that LLMs may perform better with visual numerical data than with textual numerical data. Overall, our work offers insights into the strengths and limitations of LLMs in dynamic numerical reasoning tasks, particularly in the context of geometry vs. algebra, and sets the foundation for future research on improving their performance in real-world, interdisciplinary challenges.
@@
This study evaluates LLMs within a virtual stock trading environment, focusing on their performance in visual geometric reasoning. While this controlled setting limits the generalizability to other domains, it provides valuable insights into LLMs’ capabilities. The reliance on high-quality visualizations, reflection modules, and substantial computational resources may restrict applicability in resource-constrained environments. Future research can address these limitations by broadening the scope to include diverse reasoning tasks, optimizing computational requirements, and exploring alternative modalities for more robust and generalizable assessments.

### v2

@@ 9 Ivy 3 100,000 Marketing Manager
Table 1: Agent Characteristics. Duration denotes each agent’s investment horizon, and Initial Capital indicates the starting funds for trading.
ID Ticker DPS Closing Prices Initial Quantity
1 A 22 454.17, …, 445.60 1,200
2 B 23 354.17, …, 465.80 1,000
3 C 25 500.47, …, 440.60 1,600
Table 2: Stock Details. DPS denotes dividend per share; Closing Prices are historical end-of-day closing prices; and Initial Quantity specifies the starting number of shares.
@@
We configure the Arena with varying numbers of agents and stocks, all initialized with the same capital (see Tables˜1 and 2). To validate our agents, we simulate portfolio investments on subsets of NASDAQ and CSI datasets. For NASDAQ, we run a two-month simulation (September 3–October 29, 2024) with an initial capital of 100,000 units, excluding weekends and holidays. For CSI, we conduct a one-year simulation (January 2–December 31, 2024) with the same capital, also excluding non-trading days. Both datasets, collected from Yahoo Finance and Baostock, provide daily records of opening, closing, high, and low prices, trading volume, and associated technical indicators, offering a holistic and up-to-date depiction of market behavior. Additional details are given in Appendix˜B.
Baselines.
We compare ArenaTrader against the following strategies:
1) Buy & Hold: A passive strategy that purchases assets initially and holds them throughout the period without further trading.
2) SMA (Gencay, 1996): A moving average crossover strategy used to detect bullish and bearish momentum. We evaluate window sizes [5, 10, 15, 20] and select the best based on validation results.
3) ZMR (Poterba and Summers, 1988): A zone-based reversal strategy that trades on deviations from preset bounds, with thresholds and holding periods tuned via validation.
4) MACD (Chong and Ng, 2008): A momentum strategy based on MACD and signal line crossovers, where MACD is the 12-day EMA minus the 26-day EMA, and the signal line is the 9-day EMA.
5) StockFormer (Gao et al., 2023): A hybrid model combining predictive coding and reinforcement learning to capture future dynamics and asset correlations. We follow the default setup from the original implementation.
6) TimesNet (Wu et al., 2023): Transforms 1D time series into 2D representations to model multi-period temporal variations. We follow the original setup and select the look-back window from [5, 10, 15, 20] based on validation.
7) StockMixer (Fan and Shen, 2024): A lightweight MLP that mixes indicators, time, and stocks for efficient modeling of market dynamics. We follow the settings specified in the original work.
Metrics.
We evaluate using five metrics:
1) Total Return (TR): $(C_{1}-C_{0})/C_{0}$, where $C_{0}$ and $C_{1}$ are initial and final capital.
2) Win Rate (WR): $N_{w}/N_{t}$, the ratio of profitable trading days to total trading days.
@@
3) Sharpe Ratio (SR): $(R_{p}-R_{f})/\sigma_{p}$, where $R_{p}$ is the mean daily return, $\sigma_{p}$ is the standard deviation of daily returns, and $R_{f}$ is set to 0 to emphasize risk-adjusted returns.
4) Mean Daily Return (Mean): Average daily return over the simulation.
@@
5) Return Volatility (Std): Standard deviation of daily returns, measuring risk.
5.2 Experimental Results
In the Agent Trading Arena, we configure multiple agents and LLM backbones to promote diversity. The primary agent, ArenaTrader, consistently uses our full-featured design, including visual perception and reflection, enabling focused analysis. We evaluate ArenaTrader on real-world datasets, with corresponding prompts detailed in Appendix˜C.
Model Avg. Trend TR (%) $\uparrow$ Mean (%) $\uparrow$ Std (%) $\downarrow$ WR (%) $\uparrow$ SR $\uparrow$ $\Delta$ (%) $\uparrow$
ArenaTrader w/ Qwen-VL-128k -0.4632 33.3018 2.6526 1.0150 100.00 2.6135 33.7650
ArenaTrader w/ GPT-4o-mini 5.4722 38.6460 3.0249 1.4235 100.00 2.1249 33.1738
ArenaTrader w/ Gemini-1.5 7.2743 41.3264 3.1946 0.1588 100.00 20.1128 34.0521
ArenaTrader w/ Qwen-VL-32k -0.5216 35.8150 3.1213 1.6079 90.00 1.9412 36.3366
ArenaTrader w/ GPT-4o 3.5881 47.6851 3.6095 0.5327 100.00 6.7765 44.0970
Table 3: Investment Performance of ArenaTrader with Different LLMs. Avg. Trend denotes the average market trend, and $\Delta$ indicates the gain relative to this trend. GPT-4o achieves the highest returns and stability, demonstrating superior data-driven reasoning among vision-capable LLMs. Bold and underlined values denote the best and second-best results, respectively.
Refer to caption Refer to caption
(a) Gemini-1.5 (b) Qwen-VL-128k
Refer to caption Refer to caption
(c) GPT-4o-mini (d) GPT-4o
Figure 5: ArenaTrader Performance vs. Market Volatility. Total return of ArenaTrader in the Agent Trading Arena, compared with the most volatile single stock and the average return across all stocks. This comparison highlights the agent’s robustness under varying market conditions.
Vision-Enhanced Agent Performance.
Tables˜3 and 5 report the returns of ArenaTrader under different vision-capable LLMs. GPT-4o (Hurst et al., 2024) achieves the highest TR, followed by Gemini-1.5 (Reid et al., 2024). For volatility-adjusted returns, GPT-4o again leads, with Qwen-VL-32k (Wang et al., 2024b) second. These results highlight GPT-4o’s strong data-driven reasoning and robustness across varying volatility regimes, underscoring its ability to integrate visual and numerical signals in dynamic markets.
Strategy NASDAQ CSI
TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$ TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
Buy & Hold 6.06 0.155 0.916 61.54 0.169 0.73 0.004 0.403 43.15 0.010
SMA 0.19 0.006 0.561 46.15 0.011 -3.45 -0.014 0.319 33.20 -0.044
ZMR 1.04 0.030 0.814 41.03 0.037 1.24 0.006 0.361 41.91 0.016
MACD 5.62 0.146 1.033 58.97 0.141 -0.67 -0.002 0.373 40.66 -0.006
StockFormer (Gao et al., 2023) 11.23 0.281 0.121 56.41 0.231 14.83 0.636 1.128 41.08 0.056
TimesNet (Wu et al., 2023) 11.41 0.283 1.035 56.41 0.273 32.93 0.139 2.053 48.55 0.068
StockMixer (Fan and Shen, 2024) 14.64 0.357 1.173 56.41 0.305 43.00 0.153 0.907 27.80 0.168
ArenaTrader w/ GPT-4o (Ours) 15.99 0.450 1.110 58.97 0.348 48.93 0.176 1.432 51.45 0.123
Table 4: Strategy Performance on NASDAQ and CSI. Our method demonstrates strong robustness and generalization, achieving near-optimal performance without additional training.
Evaluation on Real-World Data.
Using GPT-4o as the backbone, Table˜4 presents results on NASDAQ and CSI. Training details for StockFormer (Gao et al., 2023), TimesNet (Wu et al., 2023), and StockMixer (Fan and Shen, 2024) are provided in Appendix˜B. Despite extensive training, these models underperform our method, which uses a 10-day input window selected via ablation. The Sharpe Ratio (SR) of NASDAQ-100 benchmark during the same period is 0.189, while CSI 300 records 0.054. In contrast, our model achieves 0.348 on NASDAQ-100 (an improvement of 0.159) and 0.123 on CSI 300 (an improvement of 0.069).
(a) NASDAQ
(b) CSI
Figure 6: Backtesting Results on Real-World Stock Datasets. Two-month backtest on NASDAQ and one-year backtest on CSI. The proposed method consistently delivers near-optimal performance and outperforms all baselines, particularly during periods of high market volatility.
Building on these results, Fig.˜6 provides a detailed analysis of portfolio trajectories and market conditions, illustrating our agent’s near-optimal performance, especially during high-volatility periods. Notably, our approach operates without task-specific retraining, relying solely on historical price data. This design underscores the robustness and strong generalization ability of our agent in real-world, dynamic market environments.
5.3 Ablation Studies
We conduct ablation studies in the Arena to evaluate the impact of input formats and the reflection mechanism on agent performance. To further assess the effect of sequence length across textual and visual modalities, we perform controlled backtesting on NASDAQ. Backtesting mitigates dynamic data fluctuations, allowing sequence length to be isolated as the primary variable.
Model Textual Visual TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
w/o Reflection
LLaMa-3 (Dubey et al., 2024) 4.09 0.686 2.270 62.50 0.196
DeepSeek (Liu et al., 2024a) 27.31 2.248 2.755 92.21 1.530
Qwen-2.5 (Yang et al., 2024) 30.37 2.680 1.459 93.75 1.703
Gemini-1.5 (Reid et al., 2024) 14.32 1.221 2.778 83.84 1.729
19.04 1.604 1.165 90.91 1.376
23.96 1.981 1.334 100.00 1.485
GPT-4o (Hurst et al., 2024) 13.04 1.139 1.929 54.55 0.590
17.07 1.454 1.504 72.73 0.967
26.18 2.157 2.058 90.91 1.048
w/ Reflection
LLaMa-3 (Dubey et al., 2024) 10.25 1.667 2.403 66.67 0.694
DeepSeek (Liu et al., 2024a) 30.62 2.470 1.512 100.00 1.633
Qwen-2.5 (Yang et al., 2024) 38.91 3.324 1.073 100.00 3.099
Gemini-1.5 (Reid et al., 2024) 29.45 2.519 3.012 93.51 2.210
37.01 2.911 1.235 100.00 2.357
41.33 3.195 0.159 100.00 20.113
GPT-4o (Hurst et al., 2024) 33.65 2.671 2.171 98.70 2.142
35.76 2.821 0.678 100.00 4.159
47.69 3.610 0.533 100.00 6.777
Table 5: Performance Comparison with and without Reflection Using Textual and Visual Inputs. Agents using visual data consistently outperform those using text alone, while combining both modalities yields the strongest results.
Textual vs. Visual Inputs.
We compare agents using textual-only inputs with those leveraging visualizations (scatter plots, line charts, bar graphs). Non-vision LLMs receive text only without reflection, while vision-capable LLMs are evaluated under text-only, visual-only, and combined-input settings. Corresponding prompts are provided in Appendix˜C, and details of LLM choices are in Appendix˜D. Results in Table˜5 show that visual inputs substantially improve returns, with combined inputs achieving the best overall performance.
Impact of Reflection.
Incorporating the reflection module further enhances performance across all LLMs. For GPT-4o, TR improves from 33.65% with text-only to 47.70% with visual input, a relative gain of 41.7%. Beyond TR, reflection also boosts SR and WR, indicating higher profitability, stability, and consistency in trading outcomes. These findings confirm that strategic reflection, particularly with visual data, strengthens agents’ reasoning and decision-making under volatile market conditions.
Input Window TR $\uparrow$ Mean $\uparrow$ Std $\downarrow$ WR $\uparrow$ SR $\uparrow$
Textual 5 days 3.63 0.102 1.588 62.5 0.064
10 days 5.58 0.145 1.334 60.0 0.109
15 days 6.47 0.171 1.681 57.5 0.102
20 days 1.67 0.050 1.296 62.5 0.039
Visual 5 days 8.21 0.207 1.396 57.5 0.149
10 days 6.90 0.174 1.164 67.5 0.149
15 days 6.90 0.174 1.216 60.0 0.143
20 days 2.81 0.075 1.101 62.5 0.068
Textual
Visual
5 days 14.33 0.372 1.719 60.0 0.216
10 days 15.99 0.450 1.110 59.0 0.348
15 days 14.73 0.380 1.684 57.5 0.226
20 days 13.51 0.352 1.577 65.0 0.223
Table 6: Performance with Textual and Visual Inputs Across Different Window Sizes. Window denotes the number of consecutive trading days provided as historical input before the current trading day. Textual + Visual consistently achieves the best performance, particularly with a 10-day window, highlighting the importance of both modality choice and time-window design.
Impact of Time-Series Length.
Experiments with GPT-4o as the backbone are summarized in Table˜6, which reports performance across modalities and time windows. Overall, “Textual + Visual” consistently outperforms single modalities, particularly with a 10-day window, achieving the highest TR (15.99%), SR (0.348), and Mean (0.450%). Visual inputs perform best in the 5–10 day range, offering strong returns, stability, and win rates. Performance declines at 20 days, likely due to noise accumulation or overfitting. These results underscore the importance of both multi-modal fusion and careful time-window selection for robust time-series decision-making.
@@ 6 Conclusion
We presented the Agent Trading Arena, a closed-loop, zero-sum simulation framework designed to model complex market dynamics and evaluate LLMs on numerical reasoning tasks. Our experiments show that while LLMs struggle with plain-text numerical data, they achieve significantly higher returns when supported by visual representations, underscoring the importance of charts and graphs in complex decision-making. Integrating a reflection module further improves performance by enabling iterative analysis and strategic refinement. Validation on both NASDAQ and CSI confirms that vision-enhanced LLM agents consistently outperform text-only counterparts without additional training. Overall, this work highlights both the capabilities and limitations of LLMs in numerical reasoning and establishes the Agent Trading Arena as a robust testbed for advancing research in adaptive, real-world financial applications.
@@
This study investigates the visual reasoning abilities of LLMs in a virtual stock trading environment. While such a controlled setup enables systematic analysis, it inevitably simplifies many aspects of real-world financial systems, including uncertainty, feedback loops, and human interactions. Bridging the gap between virtual environments and real-world applications therefore remains a significant challenge.
A promising direction for future work is to explore how agents can be gradually integrated into real financial workflows, interacting with live data, adapting to real-time signals, and contributing meaningful decisions or strategies. This would move beyond passive evaluation toward active participation, where agents not only respond to environments but also generate novel behavioral data and insights to inform downstream analysis, model training, and system design. Expanding the task scope and incorporating more diverse modalities may further support the development of generalizable and robust evaluation frameworks.